# Rainbow chains and numerical renormalisation group for accurate chiral conformal spectra

### v1

@@ 4.1 Example: the three-state Potts model
Finally, we consider the ground-state entanglement spectrum of (10) in the presence of strong boundary fields $h_{L,R}$. Upon the conformal mapping (2), this corresponds to a finite-temperature BCFT with free boundary conditions on the side of the entanglement cut (where the Hamiltonian was not changed) and fixed ones on the other side. The spectrum for this setup was predicted [38, 16] to contain the conformal towers $(1/8)\oplus(13/8)$ and $(1/40)\oplus(21/40)$ in the limit of strong positive (negative) boundary fields, even though these do not appear in the spectrum of the boundaryless critical Potts model [43, 44]. We indeed recover these in our simulations, see Fig. 5(d,e); the two conformal towers are readily distinguished by their $\mathbb{Z}_{2}$ symmetry eigenvalues.
The deviation of entanglement energies from their thermodynamic-limit value scales with the effective system size $w$ as $1/w^{2}$, similar to the free-fermion systems in Sec. 3 [Fig. 5(f)]. This indicates that MPS with modest bond dimensions are able to accurately represent the leading Schmidt states even for these large systems, without limitations of a finite correlation length. Empirically, we find that the deviation depends on the MPS bond dimension $\chi$ in this limit as $1/\chi^{2}$ [Fig. 5(g)].
@@ 5 Conclusion
In this work, we have developed a method to obtain detailed spectral information about conformal field theories using entanglement spectra. By exploiting the conformal mapping between reduced and thermal density matrices [14], we designed chain Hamiltonians whose entanglement spectra match the energy spectrum of long conformal chains, thereby removing the obstacle of logarithmic finite-size scaling [13, 15]. The interaction terms in the bulk of such a chain decay exponentially, similar to the so-called rainbow chains [19, 20, 21, 22]. We first demonstrated our method on free-fermion Hamiltonians, where we could obtain entanglement spectra from single-particle calculations [26]. Indeed, we found that entanglement spectra match those of the expected chiral free-fermion and Ising CFTs, with up to 11 levels of the conformal tower clearly resolved in each symmetry sector: This is a significant improvement over uniform chains and the edge spectra of narrow cylinders, where the entanglement conformal spectra can often only resolve the first few states.
@@
We also proposed a Wilsonian numerical renormalisation group (NRG) algorithm for obtaining low-energy eigenstates for rainbow chains that represent interacting (non-free-fermion) Hamiltonians. In contrast to earlier work [39], we find that this algorithm clearly outperforms such alternatives as DMRG on the problem at hand; most importantly, it can handle chains with an arbitrarily wide dynamical range of Hamiltonian terms, as they are optimised hierarchically. Furthermore, the wave function is returned as a folded MPS [Fig. 4(b)], which is well-suited to the volume-law entanglement structure of the rainbow ground state. We benchmarked the NRG algorithm on the three-state Potts model: By computing ground- and excited-state entanglement spectra in different boundary conditions, we were able to obtain the spectrum of every conformal tower of the chiral $\mathcal{M}_{5}$ minimal model with excellent accuracy and level of detail.
As it makes no strong assumptions about the structure of the underlying Hamiltonian, our NRG approach is in principle applicable to any conformal quantum critical point in $(1+1)$ dimensions. This will allow us to identify the underlying CFTs through access to their full spectra rather than only, e.g., their central charge. For such applications, it will be important to understand the structure of the NRG flow near the critical point so it can be distinguished from the fixed points corresponding to the surrounding phases.
@@
The structure of the NRG fixed-point tensor itself will also be an interesting target for exploration, as it provides a compact representation of every conformal tower of the underlying CFT. Indeed, these can all be extracted from the same MPS, with tensors converging to the fixed-point one, by changing only the final tensor as boundary conditions are changed (cf. Fig. 5). Furthermore, the NRG procedure has a clear sense of directionality, so an MPS made of fixed-point tensors with periodic boundary conditions may well encode the density matrix of a genuinely chiral CFT, similar to the boundary MPS that represent the reduced density matrix of a two-dimensional PEPS state [6]. Comparing these two approaches to obtain approximate MPS representations of chiral conformal spectra may also provide new insights into the representability of chiral topological order with PEPS [11].
@@
Even more importantly, computing entanglement spectra only requires wave functions, not necessarily Hamiltonians, at the critical point. This will allow us to analyse critical wave functions obtained by other means and extract its underlying CFT or even parent Hamiltonian. This is a particularly promising feature for capturing edge theories of topological states (e.g., chiral spin liquids) constructed through a parton construction. Topological transitions between such phases would normally be detected through entanglement properties [45, 46, 47]; we expect that detailed access to the evolution of the entanglement spectrum across the transition will yield new insights into their mechanism.
@@ Acknowledgements
Appendix A Effective conformal system size
@@
In the first approximation, the conformal transformations (2, 4) map reduced density matrices of the rainbow-deformed chains/rings to the thermal density matrix of a system of size $(L/2-1)\Delta$ (cf. Fig. 1). This picture, however, is not completely accurate for the chains we really use in the simulations, mostly because the effective size of the segment of the chain across the entanglement cut is scaled by a factor of $\Delta/2$ to match the order of magnitude of its neighbours. [^4] As an extreme example, at $\Delta=0$, all $w_{i}=0$, so every site should be mapped onto $z_{i}=\pm 1$, leading to a zero term in the Hamiltonian across the entanglement cut. Instead, the rainbow-deformed chain construction described in Sec. 2 reduces to a uniform chain.
A better estimate of the effective system size $w$ can be obtained from assigning a segment of length $1/J_{i}$ to every hopping term (cf. Sec. 2) and scaling the resulting chain to the segment $z\in(-\pi/4,\pi/4)$ that appears in the conformal map (2) (or the circle $\theta\in[0,2\pi)$ for rings). This yields
$w_{1}$ $=-\log\tan\left(\frac{\pi}{4}\frac{1/J_{\mathrm{middle}}}{\sum_{i}1/J_{i}}\right)$ (A.1a)
$w_{2}$ $=-2\log\tan\left(\pi\frac{1/J_{\mathrm{middle}}}{\sum_{i}1/J_{i}}\right)$ (A.1b)
for chains and rings, respectively, where $J_{\mathrm{middle}}$ is the coupling strength of the bond(s) across the entanglement cut. In the limit $L\Delta\gg 1$, both expressions reduce to
$w\to\frac{L\Delta}{2}+f(\Delta);$ (A.2)
the function $f(\Delta)$ can only be obtained analytically for the rainbow chain of Fig. 1(c), where the sums in (A.1) are over geometric series:
@@
$f_{\mathrm{rainbow}}(\Delta)=-\log\left[\frac{\pi}{8}(e^{\Delta}-1)\right].$ (A.3)
$f(\Delta)$ is tabulated for the values of $\Delta$ that appear in the main text in Table 2; convergence to the asymptotic form (A.2) appears to be exponential in $L\Delta$.
$\Delta$ chain (3) rainbow chain ring (5)
$1/64$ 4.8364 5.0858 8.3021
$1/32$ 4.1276 4.3848 6.9002
$1/16$ 3.4032 3.6759 5.4827
$1/8$ 2.6476 2.9510 4.0339
$1/4$ 1.8294 2.1934 2.5226
$1/2$ 0.8863 1.3675 0.8863
Table 2: Deviation of conformal chain length $w_{1,2}$ from $L\Delta/2$ for the three types of deformed chain and the values of $\Delta$ considered in this work in the limit of $L\Delta\gg 1$.
We can extend the same argument to Hamiltonians with single- and two-site terms, treating both kinds of term as a segment – this is particularly natural for the Majorana representation (8c) of the TFI chain. Eqs. (A.1, A.2) remain applicable, although $\Delta$ must be replaced with $\Delta/2$, the effective length of these “segments.”

### v2

@@ 4.1 Example: the three-state Potts model
Finally, we consider the ground-state entanglement spectrum of (18) in the presence of strong boundary fields $h_{L,R}$. Upon the conformal mapping (2), this corresponds to a finite-temperature BCFT with free boundary conditions on the side of the entanglement cut (where the Hamiltonian was not changed) and fixed ones on the other side. The spectrum for this setup was predicted [39, 16] to contain the conformal blocks $(1/8)\oplus(13/8)$ and $(1/40)\oplus(21/40)$ in the limit of strong positive and negative boundary fields, respectively, even though these do not appear in the spectrum of the boundaryless critical Potts model [44, 45]. We indeed recover these in our simulations, see Fig. 7(d,e); the two blocks are readily distinguished by their $\mathbb{Z}_{2}$ symmetry eigenvalues.
4.2 Parameter dependence
Finite-$\chi_{\mathrm{NRG}}$ effects on NRG stability.
Figure 8: (a) Evolution of the lowest 10 NRG eigenstates in each $\mathbb{Z}_{3}$ charge sector for $\Delta=1/4$ and $\chi_{\mathrm{NRG}}=400$ (top) and 800 (bottom), diverging to the paramagnetic and ferromagnetic fixed points, respectively.
(b) NRG spectrum at the final iteration of the two trajectories. For $\chi_{\mathrm{NRG}}=800$, the ground state becomes degenerate across the three charge sectors, indicating spontaneous symmetry breaking.
(c) Evolution of the NRG ground state for three values of $\Delta$ and several values of $\chi_{\mathrm{NRG}}$. Increasing the bond dimension keeps the critical fixed point stable for longer, up to some value of $\chi_{\mathrm{NRG}}$. Achieving the same degree of stability requires substantially higher $\chi_{\mathrm{NRG}}$ for smaller values of $\Delta$.
[…]
Keeping a limited number $\chi_{\mathrm{NRG}}$ of lowest-energy states in NRG (that is, constructing an MPS representation with bond dimension capped at $\chi_{\mathrm{NRG}}$) perturbs the ideal RG flow by truncating the overlap of the ground state with higher-energy eigenstates of shorter segments. As shown in Fig. 8(a), this is not a relevant perturbation: the critical spectrum in Fig. 7(a) is stable for a large span of NRG steps. However, as the critical point is an unstable RG fixed point, these perturbations are magnified over time and eventually drive the flow to one of the two stable fixed points, corresponding to a noninteracting paramagnet ($J\to 0$) and a nondynamical Potts ferromagnet ($h\to 0$), the spectra of which are shown in Fig. 8(b). The significance of this effect also depends on $\Delta$, which controls the separation of scales between NRG steps: For large $\Delta$, high-energy states have a smaller effect on subsequent steps, so a smaller $\chi_{\mathrm{NRG}}$ keeps the NRG procedure stable for more steps, as illustrated in Fig. 8(c).
[…]
Finite-$\chi$ effects on the unzipping algorithm.
From here on, we consider the NRG wave function obtained for $\Delta=1/4,\chi_{\mathrm{NRG}}=400$ at the renormalised critical point $g_{c}$, which we take to be a faithful representation of the three-state Potts CFT wave function on the rainbow chain. However, to extract such quantities as the von Neumann entropy and the entanglement spectrum, the folded MPS representation needs to be unzipped into a regular MPS using the algorithm in Fig. 6(f–h), which is controlled by a new hyperparameter, the bond dimension $\chi_{\mathrm{(unzip)}}$ of this new MPS.
Figure 9: (a) Ground-state entanglement entropy of the critical three-state Potts model as a function of system size, obtained using the unzipping algorithm of Fig. 6(f–h) with MPS bond dimension $\chi=400$ (red) and 800 (blue). They obey the volume law with $\ell(\Delta)$ taken from the transverse-field Ising model (black line) up until saturation due to finite $\chi$ (dashed lines). The saturated values are plotted as a function of $\chi$ (green): they fit well to $S_{\mathrm{sat}}=\log\chi-\mathrm{const.}$ (green line).
(b) Entanglement gap as a function of effective system size $L_{\mathrm{eff}}=\frac{12}{c}S_{\mathrm{vN}}$. They follow (13) (black line, quadratic correction in dashes) up to saturation, where $\delta_{\mathrm{ent}}$ falls below its expected magnitude.
(c) Entanglement spectrum as a function of system size obtained with $\chi=400$ (grey) and 800 (black). Below saturation, they convergence towards the expected thermodynamic limit, with little $\chi$-dependence; above it, they diverge again, more strongly for smaller $\chi$. Fig. 7 uses $L=256$ (blue symbols), where the multiplets are the narrowest.
In Fig. 9(a), we show the von Neumann entropy obtained for a range of different system sizes and values of $\chi$. The entropy of the exact wave function is expected to increase according to the volume law $S\simeq\frac{c}{12}L\Delta$ [cf. (11)]; however, this is capped by the MPS representation, which can only capture $S\leq\log\chi$. The numerical results are consistent with this picture: Before saturation, the correction $\ell(\Delta)$ to the effective chain length matches exactly that for the TFIM (9), highlighting its geometrical origin. The saturation value of $S_{\mathrm{vN}}$, however, differs from $\log\chi$ by a constant offset, indicating some surviving structure even in this limit.
Before saturation, the entanglement gap $\delta_{\mathrm{ent}}$ (defined as half the gap between the lowest zero-charge CFT levels, $n=0,2$) also follows the scaling (13) found previously for free-fermion systems [Fig. 9(b)]. However, upon saturation, they deviate downwards from this scaling. This is to be expected, since the entanglement entropy represented by the MPS is maximised if all Schmidt vectors are equal, that is, the gap vanishes: As the MPS truncation algorithm tries to preserve as much of the entanglement entropy as possible, it pushes towards such a result.
Finally, we plot the rescaled entanglement spectrum in the zero-charge sector in Fig. 9(c). Before the saturation of entanglement entropy, the dependence of this spectrum on the bond dimension is negligible, and the spread of each conformal multiplet decreases as the system size increases, presumably with the same power laws as in the free-fermion case. By contrast, as the entanglement entropy saturates, the multiplets start to broaden again; the extent of this broadening strongly depends on the bond dimension. Our attempt to extract the relevant scaling laws proved fruitless, as the $\chi$-dependence of the spectrum for computationally accessible values shows many discrete jumps on top of any such scaling limit.
@@ 5 Conclusion
In this work, we have developed a method to obtain detailed spectral information about conformal field theories using entanglement spectra. By exploiting the conformal mapping between reduced and thermal density matrices [14], we designed chain Hamiltonians whose entanglement spectra match the energy spectrum of long conformal chains, thereby removing the obstacle of logarithmic finite-size scaling [13, 15]. The interaction terms in the bulk of such a chain decay exponentially, similar to the so-called rainbow chains [19, 20, 21, 22]. We first demonstrated our method on free-fermion Hamiltonians, where we could obtain entanglement spectra from single-particle calculations [26]. Indeed, we found that entanglement spectra match those of the expected chiral free-fermion and Ising CFTs, with up to 11 levels of conformal blocks clearly resolved in each symmetry sector: This is a significant improvement over uniform chains and the edge spectra of narrow cylinders, where the conformal entanglement spectra can often only resolve the first few states.
@@
We also proposed a Wilsonian numerical renormalisation group (NRG) algorithm for obtaining low-energy eigenstates for rainbow chains that represent interacting (non-free-fermion) Hamiltonians. In contrast to earlier work [40], we find that this algorithm clearly outperforms such alternatives as DMRG on the problem at hand; most importantly, it can handle chains with an extremely wide dynamical range of Hamiltonian terms, as they are optimised hierarchically. Furthermore, the wave function is returned as a folded MPS [Fig. 6(b)], which is well-suited to the volume-law entanglement structure of the rainbow ground state. We benchmarked the NRG algorithm on the three-state Potts model: By computing ground- and excited-state entanglement spectra for different boundary conditions, we were able to obtain the spectrum of every conformal tower of the chiral $\mathcal{M}_{5}$ minimal model with excellent accuracy and detail.
As it makes no strong assumptions about the structure of the underlying Hamiltonian, our NRG approach is in principle applicable to any conformal quantum critical point in $(1+1)$ dimensions. This allows us to identify the underlying CFTs through access to their full spectra, a major improvement over the entanglement entropy, which only recovers the central charge. In future work, it will be interesting to explore how further properties of the CFT (e.g., the list of primary fields, their spins and fusion rules) can be extracted from wave functions or NRG fixed-point tensors: Since wave functions in different symmetry sectors and boundary conditions can be extracted from the same NRG flow, only changing the final tensor in the MPS, all this CFT information must be encoded directly in the NRG flow and in the fixed-point tensor in particular.
@@
Furthermore, the NRG procedure has a clear sense of directionality, so an MPS made of fixed-point tensors with periodic boundary conditions may well encode the density matrix of a genuinely chiral CFT, similar to the boundary MPS that represent the reduced density matrix of a two-dimensional PEPS state [6]. Comparing these two approaches to obtain approximate MPS representations of chiral conformal spectra may also provide new insights into the representability of chiral topological order with PEPS [11].
@@
Even more importantly, computing entanglement spectra only requires wave functions, not necessarily Hamiltonians, at the critical point. This will allow us to analyse critical wave functions obtained by other means and extract its underlying CFT or even parent Hamiltonian. This is a particularly promising feature for capturing edge theories of topological states (e.g., chiral spin liquids) obtained from a parton construction. Topological transitions between such phases would normally be detected through entanglement properties [46, 47, 48]; we expect that detailed access to the evolution of the entanglement spectrum across the transition will yield new insights into their mechanism.
@@ Acknowledgements
Appendix A Entanglement entropy in rainbow chains
The conformal transformation (2) maps the reduced density matrix of a uniform chain of length $L$ to the thermal density matrix of the CFT on an interval of length $w_{0}\sim\log L$ [Fig. 1(a)]; as a result, the von Neumann entropy of the former equals the thermal entropy of the latter. The latter is extensive in $w_{0}$; the factor of proportionality is given by the Cardy–Calabrese formula [18] as
$S=\frac{c}{6}\log L+\tilde{c}=\frac{c}{6}w_{0}+\tilde{c}^{\prime}.$ (A.1)
Rainbow chains.
As explained in Sec. 2, chains with non-uniform couplings can be represented as a discretisation of the continuum CFT with non-uniform segment lengths that scale as $\ell_{i}\sim 1/J_{i}$. Therefore, upon the conformal transformation (2), their reduced density matrices map to the thermal density matrix on a system of length
$w_{1}=-\log\tan\left(\frac{\pi}{4}\frac{1/J_{\mathrm{middle}}}{\sum_{i}1/J_{i}}\right).$ (A.2)
For the rainbow chain, the sum in the denominator is a geometric series, summing to
$w_{1,\mathrm{rainbow}}=\log\tan\left(\frac{\pi}{4}\frac{1}{1+2\sum_{i=1/2}^{(L-1)/2}e^{i\Delta}}\right)\simeq\frac{(L-1)\Delta}{2}-\log\left[\frac{\pi}{8}(e^{\Delta}-1)\right]$ (A.3)
for $L\Delta\gg 1$, whence the entanglement entropy is given by (A.1) as
$S$ $\simeq\frac{c}{12}\left[L\Delta+\ell(\Delta)\right];$ (A.4)
$\ell(\Delta)$ $=-2\log(e^{\Delta}-1)-\Delta+\mathrm{const.}\simeq-2\log\Delta+\mathrm{const.}$ (A.5)
@@ Appendix A Entanglement entropy in rainbow chains
for $\Delta\ll 1$, in accordance with Ref. [20]. This dependence on $\Delta$ comes entirely from the centre of the chain: The effective length of the bond across the entanglement cut is scaled by a factor of $\simeq\Delta/2$ to match the order of magnitude of its neighbours; [^3] As an extreme example, at $\Delta=0$, all $w_{i}=0$, so every site should be mapped onto $z_{i}=\pm 1$, leading to a zero term in the Hamiltonian across the entanglement cut. Instead, the rainbow chain reduces to a uniform one. therefore, sites after the conformal mapping are not distributed uniformly near the corresponding end of the chain.
Conformal chains.
For the conformal chain (3), we can no longer find $\ell(\Delta)$ in closed form. However, for $L\Delta\gg 1$, the effective segment lengths still grow exponentially except for the very ends of the chain: As a result, the effective system size only changes by a constant factor, so (A.4, A.5) will still hold, albeit with a different constant offset.
Rings.
Similar to (A.2), the effective system size onto which the conformal ring geometry in Fig. 1(d) is mapped by the transformation (5) is given by
$w_{2}=-2\log\tan\left(\pi\frac{1/J_{\mathrm{middle}}}{\sum_{i}1/J_{i}}\right).$ (A.6)
The sum has no closed form for the conformally deformed coupling terms (6). Similar to the previous case, however, these can be replaced with exponentially decaying ones at the cost of adding a constant offset to $w_{2}$. The resulting geometric series again yields the volume-law scaling (A.4) with
$\ell(\Delta)\simeq-4\log\Delta+\mathrm{const.};$ (A.7)
$\ell(\Delta)$ is “doubled” since now both ends of the conformally transformed chain are deformed from uniformity.
Hamiltonians with on-site terms.
As illustrated in Fig. 2, both bond and on-site terms in the Hamiltonian effectively stand for the CFT Hamiltonian integrated along a segment: This doubles the number of discretisation points and thus effectively halves $\Delta$ in the construction of conformal/rainbow chains. Therefore, we expect that (A.4) applies with $\ell(\Delta/2)$ replacing $\ell(\Delta)$, as indeed found numerically for the tight-binding and TFI models, see Fig. 5(d).
Finally, it is worth noting that $\ell(\Delta)$ only appears to depend on the details of the geometry, not of the specific Hamiltonian: Values of $\ell(\Delta)$ obtained for the TFI Hamiltonian match the entanglement entropy of the three-state Potts model accurately, as shown in Fig. 9.
A.1 Entanglement spectrum
In addition to overall chain length, the rescaling of the central bond also affects the layout of post-conformal-transformation lattice sites. For a rainbow chain with $\Delta\ll 1$ in particular, the middle of the chain becomes nearly uniform in $z$-space, so the sites are mapped to
$w_{n}$ $=\log\tan z_{n}\simeq\log(n\times\mathrm{const.})=\log n+\mathrm{const.}$ $n=\frac{1}{2},\frac{3}{2},\dots,$ (A.8)
resulting in segment lengths that scale as
$\ell_{n}$ $\approx\frac{\mathrm{d}w_{n}}{\mathrm{d}n}\simeq\frac{1}{n}$ $n$ $=1,2,\dots.$ (A.9)
As explained in Sec. 2, such non-uniformly discretised chains can be represented through Hamiltonian terms that scale as $1/\ell_{n}\simeq n$. That is, the entanglement spectrum of the rainbow chain is expected to match the energy spectrum of a chain with couplings that are uniform in the bulk but taper off linearly on one end.
Figure 10: Log-log plot of the finite-size correction to the single-particle spectrum of a tight-binding uniform chain (red), the image of a rainbow chain ($\Delta=1/8$) upon the conformal map (2) (blue), and a piecewise linear approximation of the latter (green). For the uniform chain, $\varepsilon_{\alpha}/\delta$ approaches the thermodynamic limit from below and the correction scales as $L^{-2}$; the tapered ones approach from above, scaling as $L^{-3}$ (straight lines). Inset: Profiles of the hopping term near the end of the three types of chain.
In a very long, uniform tight-binding chain near half-filling, we expect equally spaced single-particle energy levels, at energies $\alpha\delta$, where $\delta\sim 1/L$ is the single-particle gap, and $\alpha=\pm\frac{1}{2},\pm\frac{3}{2},\dots$ for an even-length chain, cf. (16). Indeed, $\varepsilon/\varepsilon_{\mathrm{min}}$ converges to odd integers. However, while the convergence for a uniform chain is quadratic, we find it to be cubic for the conformally mapped rainbow chain (Fig. 10). This is entirely due to the linear tapering of hopping terms at the end, as shown by data for an additional, piecewise linear chain.
On the flipside, for $\Delta=2$, the rainbow chain maps to a perfectly uniform chain, thus we expect that its single-particle entanglement energies (15) scale the same way as the single-particle energy levels (16) in a uniform chain. Namely, we expect that $\varepsilon_{\alpha}/\delta_{\mathrm{ent}}$ approach $\alpha$ from below and that the correction scale quadratically. In numerical tests, we verified the former prediction, but we could not access long enough chains to test the scaling.
Appendix B Entanglement entropy and gap in free-fermion chains
As argued in Sec. 3.3, the half-chain/ring entanglement spectrum matches the energy spectrum of the image of the chain/ring under the conformal maps (2)/(5), while the entanglement entropy equals the thermal entropy of the image at inverse temperature $\beta=2\pi$. For the chains and rings considered here, these images are uniformly spaced chains (except for the ends, which, however, do not affect the long-wavelength physics, cf. Appendix A.1). For $L\Delta\gg 1$, the cosine dispersions of such chains can be linearised near zero energy, resulting in fermionic modes with equally spaced single-particle energies $\varepsilon$. We define the entanglement gap $\delta_{\mathrm{ent}}$ as their spacing, since it corresponds to a unit step in the conformal tower both for tight-binding and Majorana chains. Each of these modes contributes
$\Delta S=-\left[\frac{1}{1+e^{\beta\varepsilon}}\log\left(\frac{1}{1+e^{\beta\varepsilon}}\right)+\frac{1}{1+e^{-\beta\varepsilon}}\log\left(\frac{1}{1+e^{-\beta\varepsilon}}\right)\right]=\log(1+e^{-\beta\varepsilon})+\frac{\beta\varepsilon}{1+e^{\beta\varepsilon}}$ (B.1)
to the total entropy. In the following, we drop $\beta$, as it is absorbed into our definition of the entanglement spectrum.
Tight-binding chain.
In the limit $L\Delta\gg 1$, i.e., that of small entanglement gap, we can replace the sum of (B.1) over all modes with the integral
@@ Appendix B Entanglement entropy and gap in free-fermion chains
$S\simeq\int_{-\infty}^{\infty}\frac{\mathrm{d}\varepsilon}{\delta_{\mathrm{ent}}}\left[\log(1+e^{-\varepsilon})+\frac{\varepsilon}{1+e^{\varepsilon}}\right]=\frac{\pi^{2}}{3\delta_{\mathrm{ent}}}.$ (B.2)
Majorana chain.
The situation is almost identical, except that modes with eigenvalues ±$\pm\varepsilon$ are not independent but rather creation and annihilation operators of the same Boguliubov quasiparticle. Therefore, the total entanglement entropy is
$S\simeq\int_{0}^{\infty}\frac{\mathrm{d}\varepsilon}{\delta_{\mathrm{ent}}}\left[\log(1+e^{-\varepsilon})+\frac{\varepsilon}{1+e^{\varepsilon}}\right]=\frac{\pi^{2}}{6\delta_{\mathrm{ent}}}.$ (B.3)
General Hamiltonians.
The volume-law entropy scaling (11) motivates introducing the effective chain length (12) to account for the perturbations discussed in Appendix A. In terms of this effective length, both (B.2, B.3) can be written as
$\delta_{\mathrm{ent}}\simeq\frac{4\pi^{2}}{L_{\mathrm{eff}}}.$ (B.4)
We anticipate [and indeed find for the three-state Potts model, see Fig. 9(b)] that this relation holds beyond free-fermion models: In a chain of length $L_{\mathrm{eff}}$, momentum modes are discretised in units of $2\pi/L_{\mathrm{eff}}$; due to conformal invariance between space and time, this is also the natural unit of energy. The definition of $\delta_{\mathrm{ent}}$, however, also contains a factor of $\beta=2\pi$, yielding (B.4).